Which describes the lines represented by these equations?
y=1/2x-4
2y=x+2
A. coincident lines
B. intersecting lines
C. parallel lines
D. vertical lines

Answer by stanbon(75887)

Which describes the lines represented by these equations?
y=1/2x-4
2y=x+2
A. coincident lines
B. intersecting lines
C. parallel lines
D. vertical lines
---------
The slopes are the same (m = 1/2); the y-intercepts are different.
Ans: C.
Cheers,
Stan H.